# A Novel Testing Approach for Differences Among Brain Connectomes

### v1

@@
Abstract.
Statistical analysis on non-Euclidean spaces typically relies on distances as the primary tool for constructing likelihoods. However, manifold-valued data admits richer structures in addition to Riemannian distances. We demonstrate that simple, tractable models that do not rely exclusively on distances can be constructed on the manifold of symmetric positive definite (SPD) matrices, which naturally arises in brain connectivity analysis. Specifically, we highlight the manifold-valued Mahalanobis distribution, a parametric family that extends classical multivariate concepts to the SPD manifold. We develop estimators for this distribution and establish their asymptotic properties. Building on this framework, we propose a novel ANOVA test that leverages the manifold structure to obtain a test statistic that better captures the dimensionality of the data. We theoretically demonstrate that our test achieves superior statistical power compared to distance-based Fréchet ANOVA methods.
Key words and phrases:
ANOVA, Connectomics, Riemannian geometry
1. Introduction
The analysis of data that takes values on non-Euclidean spaces has emerged as a fundamental challenge across numerous scientific disciplines, from neuroimaging to computer vision. This has driven the development of sophisticated statistical tools specifically designed for data taking values on manifolds and metric spaces.
The statistical analysis of manifold-valued data now encompasses a rich toolkit. Fréchet ANOVA (muller) extends classical analysis of variance to general metric spaces, while functional data analysis has been adapted to handle curves and shapes on manifolds (fletcher; cornea). Shape analysis methods provide powerful frameworks for comparing anatomical structures (zhao), and Riemannian techniques have proven particularly effective for analyzing brain connectivity data (logeu; LRDF06; varo).
While metric space methods offer broad generality, real-world data often possess richer structure than just distances. Many datasets naturally lie on smooth manifolds equipped with differentiable structure: brain connectivity matrices live in symmetric positive definite (SPD) manifolds, directional data lie on spheres, and shape spaces have well-understood Riemannian geometry. By exploiting this additional structure, we can develop more powerful statistical methods tailored to the specific geometry of the data.
This specialization represents a deliberate trade-off: we sacrifice the universality of metric space approaches to gain statistical power through geometric tools. Methods that leverage manifold structure can exploit tangent space approximations, parallel transport, and geodesic regression, which are absent in general metric spaces. While this may require computing gradients and other differential quantities, modern computational resources and automatic differentiation make these calculations tractable even for large-scale problems.
In this paper, we focus on developing such specialized methods for brain connectivity analysis. Brain connectomes naturally form a Riemannian manifold with well-studied geometry. The Affine Invariant Riemannian Metric (AIRM) on the SPD manifold $\mathcal{M}$ provides a principled framework for analyzing these data (logeu; LRDF06).
Statistical inference on connectome-valued data has progressed significantly. The sample Fréchet mean has been shown to possess desirable properties including existence, uniqueness for negatively-curved manifolds (karcher; kendall), consistency (batta), and asymptotic normality (kendall2; clt1). Building on these foundations, researchers have developed regression methods for manifold-valued outcomes (fletcher; cornea) and comparative analyses across populations (varo).
However, existing methods for comparing groups of connectomes are understudied. The popular Fréchet ANOVA test (muller), while applicable to general metric spaces, relies on comparing variations (sums of squared distances from data points to group means). Critically, when applied to Euclidean data, Fréchet ANOVA does not reduce to classical MANOVA tests like Wilks’ Lambda or Pillai’s trace. We propose a Riemannian generalization of classical MANOVA. Our approach leverages the differentiable structure of the SPD manifold to construct test statistics that reduce to their classical counterparts in Euclidean space while remaining computationally tractable through modern optimization techniques.
Our contributions consist of three parts. First, we introduce a test statistic that generalizes classical Wilks’ Lambda to the SPD manifold, providing a principled bridge between Euclidean and Riemannian inference. Second, we derive the asymptotic distributions of these statistics. Third, we theoretically demonstrate that our approach achieves greater statistical power than existing methods under certain assumptions.
@@ 1. Introduction
The remainder of this paper is organized as follows. In Section 2, we develop our methodological framework in three stages. We begin by introducing parameter estimation for the Mahalanobis Affine Invariant Riemannian Metric (MAIRM), a generalization of the classical affine invariant metric that relaxes sphericity assumptions. We then construct our Riemannian generalization of Wilks’ Lambda, establishing its fundamental properties and showing how it reduces to classical MANOVA in the Euclidean case. Finally, we derive the asymptotic distribution of our test statistic under the null hypothesis and analyze its behavior under alternatives, formally comparing its power properties to those of Fréchet ANOVA. Section 3 summarizes our contributions and discusses extensions for future work. Technical details on the Riemannian geometry of the SPD manifold and additional computational formulas are provided in the appendices.
2. Methods
2.1. Parameter Estimation
@@
In a number of statistical models based on Riemannian geometry ideas, it is assumed that the likelihood depends exclusively on the geodesic distance between observations and some reference point. The simplest such model is decribed in varo: for a manifold-valued random variable $\boldsymbol{C}$,
@@
$p(\boldsymbol{C})\propto\exp\left(-\frac{1}{2\sigma^{2}}\norm{\log_{\boldsymbol{C}^{\ast}}(\boldsymbol{C})}_{\boldsymbol{C}^{\ast}}^{2}\right)$
@@
$\hat{\sigma}^{2}=\frac{1}{n}\sum_{i=1}^{n}\norm{\log_{\bar{\boldsymbol{C}}_{n}}(\boldsymbol{C}_{i})}_{\bar{\boldsymbol{C}}_{n}}^{2}$
The estimators are independent.
Such a model makes a sphericity assumption. To make it explicit, consider the AIRM metric for concreteness and recall the vectorization maps (logeu):
@@
$\log_{\boldsymbol{C}}(\boldsymbol{\Lambda})$ $\mapsto\operatorname{Vect}_{I}(\boldsymbol{C}^{-1/2}\log_{\boldsymbol{C}}(\boldsymbol{\Lambda})\boldsymbol{C}^{-1/2})$
@@
However, the fact that $\bar{\boldsymbol{C}}_{n}$ converges in probability to $\boldsymbol{C}^{\ast}$ is a direct consequence of Theorem 2.3 in clt4. In addition:
Proposition 1.
@@
A sphericity assumption is also used in regression models with Euclidean predictors and connectome-valued responses. A popular model (fletcher) is based on choosing $\boldsymbol{C}^{\ast}\in\mathcal{M}$, $\{v_{k}\}_{k}\in T_{\boldsymbol{C}^{\ast}}\mathcal{M}$, $\sigma^{2}\in\mathbb{R}^{+}$ and assumming that the observations $\{(\mathbf{x}_{i},\boldsymbol{C}_{i})\}_{i}$ are distributed according to the likelihood
$p(\boldsymbol{C}_{i})\propto\exp\left(-\frac{1}{2\sigma^{2}}\norm{\log_{\boldsymbol{C}_{i}^{\ast}}(\boldsymbol{C}_{i})}_{\boldsymbol{C}_{i}^{\ast}}^{2}\right)$
@@
Existence and uniqueness of OLS estimates in general is known (fletcher). Of course, with this likelihood they are MLE estimators.
@@
(1) $p(\boldsymbol{C}_{i})\propto\exp\left(-\frac{1}{2}\operatorname{Vect}_{\boldsymbol{C}_{i}^{\ast}}(\log_{\boldsymbol{C}_{i}^{\ast}}(\boldsymbol{C}_{i}))^{T}\boldsymbol{\Gamma}^{-1}\operatorname{Vect}_{\boldsymbol{C}_{i}^{\ast}}(\log_{\boldsymbol{C}_{i}^{\ast}}(\boldsymbol{C}_{i}))\right)$
@@
Then the MLE estimators of model (1) are the OLS estimates of $\langle,\rangle^{\operatorname{Mah}}$, whose existence and uniqueness follow from (fletcher). ∎
We call the family of metrics introduced in the previous proof Mahalanobis Affine Invariant Riemannian Metric (MAIRM).
The question of estimation rises again. The procedure we propose is similar to the previous one. The $\boldsymbol{C}_{i}^{\ast}$ still satisfy
$\boldsymbol{C}_{i}^{\ast}=\mathbb{E}[\boldsymbol{C}_{i}|\mathbf{X}=\mathbf{x}]$
hence the OLS predicted values $\hat{\boldsymbol{C}}_{i}$ are consistent estimators of $\boldsymbol{C}^{\ast}_{i}$ and
(2) $\hat{\mathbf{S}}_{n}=\frac{1}{n-1}\sum_{i}\operatorname{Vect}_{\hat{\boldsymbol{C}}_{i}}(\log_{\hat{\boldsymbol{C}}_{i}}(\boldsymbol{C}_{i}))\operatorname{Vect}_{\hat{\boldsymbol{C}}_{i}}(\log_{\hat{\boldsymbol{C}}_{i}}(\boldsymbol{C}_{i}))^{T}$
is a consistent estimator of $\Gamma$.
Thus, the MAIRM and AIRM predicted values are the same, but the confidence intervals are different.
2.2. Riemannian Generalization of Wilk’s Lambda
@@
Let $\hat{C}$ denote the Fréchet sample mean of the whole sample and $\hat{C}_{l}$ denote the Fréchet sample mean of the subsample $\{C_{lj}\}_{j}$.
Proposition 3.
@@
2.3. Asymptotic Distribution of the Test Statistic
@@
Theorem 4 (Asymptotic Distribution).
Let $Y_{lj}\sim p(C)$ for $l=1,\ldots,g$ and $j=1,\ldots,n_{l}$, where
@@
where $d=p(p+1)/2$ is the dimension of the tangent space for $p\times p$ symmetric positive definite matrices.
@@
The proof proceeds by establishing the asymptotic behavior of the components of $\Lambda^{*}$ and then deriving the limiting distribution. Recall from Proposition 3 that the test statistic is defined as:
@@
We begin by applying the central limit theorem for Fréchet means (clt4). Define the chart $\phi:\text{SPD}(p)\to\mathbb{R}^{d}$ by $\phi(C)=\operatorname{Vec}_{C^{*}}(\log_{C^{*}}(C))$.
@@
Applying Theorem 2.1 of batta, we obtain:
@@
That is, the metric terms elegantly cancel, which establishes:
(5) $)\displaystyle\sqrt{n}\,\operatorname{Vec}_{C^{*}}(\log_{C^{*}}(\hat{C}$)) $\xrightarrow{d}N(0,\Gamma)$
@@
First, applying Corollary 9 with $A=\hat{C}_{l}$, $B=\hat{C}$, $C=C^{*}$, and noting that $d_{R}(\hat{C}_{l},C^{*})=O_{p}(n_{l}^{-1/2})$ and $d_{R}(\hat{C},C^{*})=O_{p}(n^{-1/2})$ from the convergence of Fréchet sample means, we obtain
@@
$u_{lj}=\operatorname{Vec}_{\hat{C}_{l}}(\log_{\hat{C}_{l}}(Y_{lj}))$
To establish the stability under change of base point, we analyze how the vectorization changes from $\hat{C}_{l}$ to $C^{*}$. Corollary 9 yields the expansion:
$\log_{\hat{C}_{l}}(Y_{lj})=\Pi_{C^{*}\to\hat{C}_{l}}\left(\log_{C^{*}}(Y_{lj})-\log_{C^{*}}(C_{l})\right)+O(n^{-1/2})$
@@
The rest of the argument proceeds exactly as in Euclidean MANOVA. ∎
2.4. Comparison to Fréchet ANOVA
The Fréchet ANOVA test (muller) provides a distribution-free alternative for testing equality of population Fréchet means and variances in general metric spaces. We now establish the precise relationship between Fréchet ANOVA and our Riemannian ANOVA test under the Riemannian normal model.
@@ 2.4. Comparison to Fréchet ANOVA
Proposition 5 (Fréchet ANOVA Non-centrality Under Normality).
@@
Suppose $Y_{lj}\sim p(C_{l}^{*})$ where the connectomes follow a Riemannian normal distribution on $\mathrm{SPD}(p)$ with population mean $C_{l}^{*}$ and covariance $\Gamma$ in the tangent space (as defined in Section 2.3).
@@
$V_{l}=\mathbb{E}[d^{2}(Y_{lj},C_{l}^{*})]=\operatorname{tr}(\Gamma)$
for all $l=1,\ldots,g$, since $\log_{C_{l}^{*}}(Y_{lj})\sim\mathcal{N}(0,\Gamma)$ in the vectorized tangent space.
@@
$V_{p}=\sum_{l=1}^{g}\lambda_{l}\mathbb{E}[d^{2}(Y_{lj},C^{*})]=\sum_{l=1}^{g}\lambda_{l}\left(V_{l}+d^{2}(C_{l}^{*},C^{*})\right),$
@@
Proposition 6 (Non-centrality under Alternative Hypothesis).
@@
    $2\log\Lambda^{*}\xrightarrow{d}\chi^{2}_{d(g-1)}(\delta)$
@@
    $\beta_{n}=P\left(-2\log\Lambda^{*}>\chi^{2}_{d(g-1),1-\alpha}\mid H_{1}\right)\to 1$
    as $n\to\infty$, establishing consistency of the test.
@@
Finally, we establish power and consistency. Since $\delta_{n}=O(n)$, the non-centrality parameter grows linearly with sample size. Therefore, for any fixed critical value $c_{\alpha}$:
$\beta_{n}=P(-2\log\Lambda^{*}>c_{\alpha}\mid H_{1})\to 1$
@@
Proposition 7 (Comparison of Non-centrality Parameters).
@@
(8) $\delta_{\text{Novel}}$ $=\frac{n\theta^{2}}{g\sigma^{2}}=O(\theta^{2})$
(9) $\delta_{\text{Fréchet}}$ $=\frac{n\theta^{4}}{g\sigma^{2}d^{2}}=O(\theta^{4})$
@@
Remark 1 (Practical Implications).
@@
3. Conclusion
We have developed a novel framework for statistical inference on brain connectivity data that leverages the full geometric structure of the symmetric positive definite manifold. Our approach addresses fundamental limitations of existing methods by introducing a Riemannian generalization of classical MANOVA that properly exploits the dimensionality of connectome data.
In the methods section, we introduced three key contributions. First, we developed the Mahalanobis Affine Invariant Riemannian Metric (MAIRM), which relaxes the sphericity assumptions inherent in existing Riemannian models while maintaining computational tractability. This framework provides consistent estimators for both location and dispersion parameters and reduces to classical multivariate analysis when applied to Euclidean data. Second, we constructed a test statistic that generalizes Wilks’ Lambda to the SPD manifold, establishing a rigorous connection between Euclidean and Riemannian inference. Our statistic respects the manifold geometry while preserving the interpretability of classical MANOVA. Third, we derived the asymptotic distribution of our test statistic under both null and alternative hypotheses, enabling practical hypothesis testing with theoretical guarantees.
The theoretical analysis revealed a striking advantage of our approach over existing methods. While Fréchet ANOVA’s non-centrality parameter scales as $O(\theta^{4})$ in the effect size, our Riemannian ANOVA achieves $O(\theta^{2})$ scaling. This quadratic versus quartic dependence translates to substantially improved power for detecting location differences, particularly for small to moderate effect sizes commonly encountered in neuroimaging studies. However, this power gain comes at the cost of stronger distributional assumptions—our method requires the data to follow a Riemannian normal distribution, whereas Fréchet ANOVA applies to general metric spaces.
Several directions for future work emerge from this foundation. First, developing goodness-of-fit tests for the Riemannian normal model would enable practitioners to assess whether our stronger assumptions are justified for their data. Second, extending our framework to handle heterogeneous covariance structures across groups would broaden its applicability. Third, incorporating covariates and regression structures while maintaining the geometric insights developed here represents an important direction for more flexible modeling. Finally, computational optimizations exploiting the exact formulas derived in our appendices could enable application to high-dimensional connectome data from modern neuroimaging studies.
The framework presented here demonstrates that carefully exploiting manifold structure can yield substantial gains in statistical efficiency. By moving beyond distance-based methods to embrace the full geometric toolkit available on Riemannian manifolds, we open new possibilities for powerful and interpretable inference on complex structured data.
Riemannian Geometry Details

### v2

@@
Abstract
The study of samples that take values in non-Euclidean metric spaces is an important part of modern statistics literature. For instance, it includes brain connectomics, an active area of neuroscience where data consists of symmetric, positive definite (SPD) matrices. Now, Fréchet ANOVA has played a significant role in the study of non-Euclidean samples, due in part to its wide applicability. This paper addresses two gaps in that literature that derive from two observations about that method. First, that it is not a generalization of MANOVA; in fact, such a generalization has not been sufficiently explored, and our first goal is to propose one. Second, that its power function has not been systematically studied; our second goal is to shed some light on that subject, including an argument for how in some scenarios our generalization is more powerful. Hopefully, these contributions will demonstrate that there are additional directions in which the study of non-Euclidean samples can go and hence spark more interest in the subject.
Riemannian geometry,
keywords:
[…]
1 Introduction
The goal of generalizing statistical methods, particularly multivariate ones, to settings in which data poses structure other than (or on top of) a regular vector is an old one. The practice of statistics requires the measure of similarity, so these generalizations are often made in the setting of metric spaces. Underlying probabilistic methods for that setting have been based on the concept of Fréchet mean (Karcher, 1977). The sample mean has been shown to possess desirable properties including existence, uniqueness for negatively curved manifolds (Kendall, 1990), consistency (Bhattacharya and Patrangenaru, 2003), and asymptotic normality (Hotz, Le and Wood, 2024). Using that concept, we can generalize traditional statistical models like linear regression to that setting as well (Thomas Fletcher, 2013).
Oftentimes, data has a layer of structure that sits on top of the metric space, that of a Riemannian manifold, allowing for additional complexity in statistical modeling, strengthening what we can say about linear regression (Cornea et al., 2017) and facilitating the construction of other models.
We obtain yet another layer of structure when we focus on specific manifolds, the one of symmetric positive definite $p\times p$ matrices, $\mathcal{M}=\mathrm{SPD}_{p}$ being the one that is of interest to us. The seminal work of several researchers (Pennec, Fillard and Ayache, 2006; Arsigny et al., 2005; Lenglet et al., 2006a) illustrate how the choice of a specific Riemannian metric turns these abstract ideas into actually computable statistical models with concrete applications to brain imaging (Varoquaux et al., 2010). This hierarchy of structure means that statistical modeling of non-Euclidean data lives on a spectrum. On the general side of the spectrum, if we make mild assumptions about the structure of the data, we obtain widely applicable models. On the specific side, if we have a concrete structure in mind, we can use additional tools.
Fréchet ANOVA (Dubey and Müller, 2019) is a seminal statistical test for comparing multiple samples of non-Euclidean data. It sits on the general side of that spectrum; it only assumes a metric space structure with mild conditions. The goal of this paper is to propose and analyze a test for a similar hypothesis on the specific side of the spectrum. The idea behind this new test is that the properties of certain Riemannian metrics on $\mathcal{M}$ are close enough to those of Euclidean space as to allow for classical ANOVA to carry over. In this way, we obtain a test that is much less general than Fréchet ANOVA, but as we show, what we lose in generality we gain in power, at least in a certain region of the parameter space.
On a more technical level, this paper addresses two big challenges. First, carrying over classical ANOVA requires us to identify a way to obtain the decomposition of squares that characterizes that method in the manifold setting. We do so by studying projections of the sample to the tangent space of the Fréchet mean. Second, we identify distributional assumptions under which we can derive an asymptotic distribution of the new test statistic. Those assumptions correspond to what we call the Riemannian normal likelihood. We also observe that said likelihood has a Riemannian metric associated to it, which we call the MAIRM.
@@ 1 Introduction
The remainder of this paper is organized as follows. In Section 2, we develop our methodological framework in three stages. We begin by introducing parameter estimation for the MAIRM, a generalization of classical affine invariant metric that relaxes sphericity assumptions. We then construct our Riemannian generalization of Wilks’ Lambda, establishing its fundamental properties and showing how it reduces to classical multivariate ANOVA in the Euclidean case. Finally, we derive the asymptotic distribution of our test statistic under the null hypothesis and analyze its asymptotic behavior under alternatives, formally comparing its power properties to those of Fréchet ANOVA. Section 3 summarizes our contributions. Technical details on the Riemannian geometry of the SPD manifold and additional computational formulas are provided in the appendix.
2 Contributions
2.1 Mahalanobis AIRM
@@
In a number of statistical models based on Riemannian geometry ideas, it is assumed that the likelihood depends exclusively on the geodesic distance between observations and some reference point. The simplest such model is decribed in Varoquaux et al. (2010): for a manifold-valued random variable $\boldsymbol{C}$,
@@
$p(\boldsymbol{C})\propto\exp\left(-\frac{1}{2\sigma^{2}}\left\lVert\log_{\boldsymbol{C}^{\ast}}(\boldsymbol{C})\right\rVert_{\boldsymbol{C}^{\ast}}^{2}\right)$
@@
$\hat{\sigma}^{2}=\frac{1}{n}\sum_{i=1}^{n}\left\lVert\log_{\bar{\boldsymbol{C}}_{n}}(\boldsymbol{C}_{i})\right\rVert_{\bar{\boldsymbol{C}}_{n}}^{2}$
These estimators are independent.
Such a model makes a sphericity assumption. To make it explicit, consider the AIRM metric for concreteness and recall the vectorization maps (Pennec, Fillard and Ayache, 2006):
@@
$\boldsymbol{\Lambda}$ $\mapsto\operatorname{Vect}_{I}(\boldsymbol{C}^{-1/2}\boldsymbol{\Lambda}\boldsymbol{C}^{-1/2})$
@@
However, the fact that $\bar{\boldsymbol{C}}_{n}$ converges in probability to $\boldsymbol{C}^{\ast}$ is a direct consequence of Theorem 2.3 in Bhattacharya and Patrangenaru (2005). In addition:
Proposition 2.1.
@@
A sphericity assumption is also used in regression models with Euclidean predictors and connectome-valued responses. A popular model (Thomas Fletcher, 2013) is based on choosing $\boldsymbol{C}^{\ast}\in\mathcal{M}$, $\{v_{k}\}_{k}\in T_{\boldsymbol{C}^{\ast}}\mathcal{M}$, $\sigma^{2}\in\mathbb{R}^{+}$ and assumming that the observations $\{(\mathbf{x}_{i},\boldsymbol{C}_{i})\}_{i}$ are distributed according to the likelihood
$p(\boldsymbol{C}_{i})\propto\exp\left(-\frac{1}{2\sigma^{2}}\left\lVert\log_{\boldsymbol{C}_{i}^{\ast}}(\boldsymbol{C}_{i})\right\rVert_{\boldsymbol{C}_{i}^{\ast}}^{2}\right)$
@@
Existence and uniqueness of OLS estimates in general is known (Thomas Fletcher, 2013). Of course, with this likelihood they are MLE estimators.
@@
$p(\boldsymbol{C}_{i})\propto\exp\left(-\frac{1}{2}\operatorname{Vect}_{\boldsymbol{C}_{i}^{\ast}}(\log_{\boldsymbol{C}_{i}^{\ast}}(\boldsymbol{C}_{i}))^{T}\boldsymbol{\Gamma}^{-1}\operatorname{Vect}_{\boldsymbol{C}_{i}^{\ast}}(\log_{\boldsymbol{C}_{i}^{\ast}}(\boldsymbol{C}_{i}))\right)$ (1)
@@
Then the MLE estimators of model (1) are the OLS estimates of $\langle,\rangle^{\operatorname{Mah}}$, whose existence and uniqueness follow from (Thomas Fletcher, 2013). ∎
We call the family of metrics introduced in the previous proof Mahalanobis Affine Invariant Riemannian Metric (MAIRM) and the likelihood above the Riemannian normal model.
2.2 Riemannian Generalization of Wilk’s Lambda
@@
Let $\hat{C}$ denote the Fréchet sample mean of the whole sample (see Appendix 1 for details) and $\hat{C}_{l}$ denote the Fréchet sample mean of the subsample $\{C_{lj}\}_{j}$.
Proposition 2.3.
@@
2.3 Asymptotic Distribution of the Test Statistic
@@
Theorem 2.4 (Asymptotic Distribution).
Let $C_{lj}\sim p(C)$ for $l=1,\ldots,g$ and $j=1,\ldots,n_{l}$, where
@@
where $d=p(p+1)/2$ is the dimension of $\mathcal{M}$.
@@
The proof proceeds by establishing the asymptotic behavior of the components of $\Lambda^{*}$ and then deriving the limiting distribution. Recall from Proposition 2.3 that the test statistic is defined as:
@@
We begin by applying the central limit theorem for Fréchet means (Bhattacharya and Patrangenaru, 2005). Define the chart $\phi:\text{SPD}(p)\to\mathbb{R}^{d}$ by $\phi(C)=\operatorname{Vec}_{C^{*}}(\log_{C^{*}}(C))$.
@@
Applying Theorem 2.1 of Bhattacharya and Patrangenaru (2003), we obtain:
@@
That is, the metric terms cancel, which establishes that:
$)\displaystyle\sqrt{n}\,\operatorname{Vec}_{C^{*}}(\log_{C^{*}}(\hat{C}$)) $\xrightarrow{d}N(0,\Gamma)$ (4)
@@
First, using Corollary A.2, with $A=\hat{C}_{l}$, $B=\hat{C}$, $C=C^{*}$, and noting that $d_{R}(\hat{C}_{l},C^{*})=O_{p}(n_{l}^{-1/2})$ and $d_{R}(\hat{C},C^{*})=O_{p}(n^{-1/2})$ from the convergence of Fréchet sample means, we obtain
@@
$u_{lj}=\operatorname{Vec}_{\hat{C}_{l}}(\log_{\hat{C}_{l}}(C_{lj}))$
To establish the stability under change of base point, we analyze how the vectorization changes from $\hat{C}_{l}$ to $C^{*}$. Using Corollary A.2, we obtain the expansion:
$\log_{\hat{C}_{l}}(C_{lj})=\Pi_{C^{*}\to\hat{C}_{l}}\left(\log_{C^{*}}(C_{lj})-\log_{C^{*}}(C_{l})\right)+O(n^{-1/2})$
@@
The rest of the argument proceeds exactly as in Euclidean MANOVA (Anderson and Olkin, 1994). ∎
2.4 Comparison to Fréchet ANOVA
The Fréchet ANOVA test (Dubey and Müller, 2019) provides a distribution-free alternative for testing equality of population Fréchet means and variances in general metric spaces. We now establish the precise relationship between Fréchet ANOVA and our Riemannian ANOVA test under the Riemannian normal model.
@@ 2.4 Comparison to Fréchet ANOVA
Proposition 2.5 (Fréchet ANOVA Non-centrality Under Normality).
@@
Suppose $C_{lj}\sim p(C_{l}^{*})$ where the connectomes follow a Riemannian normal distribution on $\mathrm{SPD}(p)$ with population mean $C_{l}^{*}$ and covariance $\Gamma$ in the tangent space (as defined in Section 2.3).
@@
$V_{l}=\mathbb{E}[d^{2}(C_{lj},C_{l}^{*})]=\operatorname{tr}(\Gamma)$
for all $l=1,\ldots,g$, since $\log_{C_{l}^{*}}(C_{lj})\sim\mathcal{N}(0,\Gamma)$ in the vectorized tangent space.
@@
$V_{p}=\sum_{l=1}^{g}\lambda_{l}\mathbb{E}[d^{2}(C_{lj},C^{*})]=\sum_{l=1}^{g}\lambda_{l}\left(V_{l}+d^{2}(C_{l}^{*},C^{*})\right),$
@@
Proposition 2.6 (Non-centrality under Alternative Hypothesis).
@@
    $n\log\Lambda^{*}\xrightarrow{d}\chi^{2}_{d(g-1)}(\delta)$
@@
    $\beta_{n}=P\left(-n\log\Lambda^{*}>\chi^{2}_{d(g-1),1-\alpha}\mid H_{1}\right)\to 1$
    as $n\to\infty$.
@@
Finally, Since $\delta_{n}=O(n)$, the non-centrality parameter grows linearly with sample size. Therefore, for any fixed critical value $c_{\alpha}$:
$\beta_{n}=P(-n\log\Lambda^{*}>c_{\alpha}\mid H_{1})\to 1$
@@
Proposition 2.7 (Comparison of Non-centrality Parameters).
@@
$\delta_{\text{Novel}}$ $=\frac{n\theta^{2}}{g\sigma^{2}}=O(\theta^{2})$ (7)
$\delta_{\text{Fréchet}}$ $=\frac{n\theta^{4}}{g\sigma^{2}d^{2}}=O(\theta^{4})$ (8)
@@
Remark 2.8 (Practical Implications).
@@
3 Conclusion
In this paper, we have shown that while Fréchet ANOVA is a widely applicable test, we can sacrifice some of its generality to obtain a test that in certain situations can be more powerful. We have proposed one such test by leveraging properties of certain Riemannian metrics on the SPD manifold to carry over the sum of squares decomposition at the heart of classical multivariate ANOVA. We have also identified distributional assumptions, namely the Riemannian normal model, under which we can derive the asymptotic distribution of the novel test statistic. Substantial technical work was done in this direction, consisting of analyzing the behavior of certain Jacobi fields to bound the possible effect of curvature on the classical decomposition of observations from the Fréchet mean. In turn, we have used that asymptotic distribution to show that the novel test is more powerful than Fréchet ANOVA in certain regions of the parameter space. In this way, we have shown that the study of the statistical properties of manifold-valued samples still offers plenty of new research directions to explore, and that the geometry of the sample space endows statisticians with a varied and powerful toolkit.
Appendix A Riemannian Geometry Background
A.1 Riemannian Geometry Review